OpenStudy (anonymous):

Write this equation in slope-intercept form standard form, and point slope form m=4,b=-3

OpenStudy (lgbasallote):

do you know the slope intercept form?

Parth (parthkohli):

Slope-intercept form:\[y = mx + b \]Point slope form:\[y - y_1 = m(x - x_1) \]Can you manage the rest?